Determine the value c so that each of the following functions can serve as a probability distribution of the discrete random variable X:

(a) f(x) = c(x2 + 4), for x = 0, 1, 2, 3;


1
Expert's answer
2021-05-07T13:44:01-0400

Given that,


"f(x) = c(x^2+4)" "x=0,1,2,3"


We know,


"\sum_{x=0}^{3}f(x) = 1"


"c(4)+c(5)+c(8)+c(3) = 1"


"30c = 1"

Hence, "c = \dfrac{1}{30}"